OpenStudy (anonymous):

Find the slope of the line described by 8x = 48 + 3y

OpenStudy (bahrom7893):

8x - 48 = 3y Slope is 8/3

OpenStudy (anonymous):

In order to find the slope of a line you can use the linear equation y=mx+b where m is our slope and b is the y intercept. You were given 8x = 48 + 3y mx=b+y as your form you need to use algebra to convert it to y=mx+b 8x=48+3y Flip the equation 3y+48=8x Subtract 48 3y=8x-48 divide by 3 y=8x/3-48/3 which is of the form y=mx+b knowing that m is a real number (containing integer, rational and real numbers) we know that 8/3 is the slope of the equation.

OpenStudy (anonymous):

8x = 48 + 3y 3y = 8x - 48 ->y = 8x/3 - 16 => Slope m = 8/3